# BECoTTA: Input-dependent Online Blending of Experts for Continual Test-time Adaptation

### v2

@@ 4.3 Main Results
As shown in Tab. 3, we display the first round CGS scenario including four tasks. Even though the target domain is the same setting as CDS-Easy, the overall performances are measured higher since the accessibility of previous domains. Our BECoTTA-L achieves +5.5%p higher performance than EcoTTA with a similar number of parameters (3.16M). In this case, the input-dependent process of BECoTTA performs well in these blurry scenarios and ultimately shows +13.4%p improvement over the source model.
@@
Zero-shot Domain Generalization (DG). In addition to continual scenarios, we further demonstrate the versatility of our method through the zero-shot evaluation on four out-of-distribution tasks: BDD100k, Mapillary, CTAV, and Synthia. We compare the zero-shot performance of models using two different backbones, Deeplab v3+ and Segformer-B2. As shown in Tab. 4, our proposed method constantly outperforms strong baselines, demonstrating the competitive potential of generalization ability over unseen domains.
@@ 4.4 Analyses and Ablations
Pseudo label analysis. In  Fig. 5, we compare the generated pseudo labels after finishing the ten rounds. Our method exhibits robustness to forgetting in pseudo-label generation compared to other models. For other baselines, there is an erosion of minor labels by the effect of dominant labels (e.g. sky). However, our BECoTTA prevents such occurrences and effectively preserves pseudo labels as each round goes by, and it demonstrates significant efficacy in preserving fine-grained labels.
@@
Table 5: Ablation study for each element of the proposed BECoTTA. We measure the average IoU among 10 rounds.
Table 6: Ablation of the quality of WAD. Our WAD has versatility with diverse augmentation methods.
DD MoDE $\Theta(D;A)$ Average IoU
@@
WAD Augmentation Round 1
B-Clear A-Fog A-Night A-Snow B-Over Avg
BECoTTA (Ours) - M 43.8 68.8 34.9 57.9 49.2 50.9
      + Style-transfer 45.6 70.8 42.6 59.7 50.8 53.9
      + Transformation 45.3 70 43.2 59.5 50.7 53.7
Table 6: Ablation of the quality of WAD. Our WAD has versatility with diverse augmentation methods.
Figure 6: Detailed Analysis. We conduct the ablation study for the number of experts and $k$. Moreover, we analyze the relation between pre-defined WAD and target domains.
Ablation for each element. We observe the variation of the main components: Domain Discriminator (DD), MoDE, and domain-expert synergy loss $\Theta(D;A)$. As shown in Tab. 6, relying solely on DD results in only +0.04%p improvement compared with the source only. However, when incorporating the MoDE and $\Theta(D;A)$, there are +5.36%p and +6.12%p improvement respectively, comparing with the source model.
Quality of WAD. We adopt different augmentation methods to build realistic WAD. We utilize TSIT (Jiang et al., 2020) for style-transfer and PyTorch transformation (e.g., ColorJitter, RandomGrayscale), same as EcoTTA (Song et al., 2023). As shown in Tab. 6, we verify that the quality of WAD is a less important factor to have an effect on our domain-adaptive architecture. This demonstrates that our BECoTTA is implemented with various augmentations, showing its potential for expansion in diverse situations.
Effect of the number of $\bm{N}$ and $\bm{K}$. Our BECoTTA allows for the adjustment of the number of parameters and the complexity by varying $N$ (the whole number of experts) and $K$ (the number of selected experts). In Fig. 6 left, we illustrate these effects on the CDS-Hard scenario. From this, it is evident that the optimal results are obtained with $N$=6 and $k$=3. However, through numerous experiments, we find that regardless of the $N$ and $K$ chosen, our method consistently outperforms the existing baselines, demonstrating that our method is robust to hyperparameter settings.
Relation between WAD and $\bm{X_{t}}$. During the CTTA, we generate pseudo domain labels $d$ for each $X_{t}$ (target domains) through DD. These labels are used for initial domain-router assignment to facilitate domain-adaptive routing. Therefore, the relevance between the WAD and the target domain is crucial. In  Fig. 6 right, we represent the relationship between pre-defined WAD and $X_{t}$, and our DD faithfully reflects this connection.
@@ Appendix A Baselines
In this section, we provide the details of the TTA baselines we use in our main paper.
@@ B.1 Implementation Details
Dataset WAD Target domains
@@ B.2 The details of BECoTTA
Details of w/ & w/o WAD. To ensure a fair comparison, we conduct all experiments in w/ & w/o WAD settings. In the case w/o WAD, it is impossible to collect priors for domain candidates, therefore we adopt a single router $G$, similar to the conventional stochastic routing (Zuo et al., 2021). In settings w/ WAD, as described in the paper, we employ domain-adaptive routing and a domain-experts synergy loss. This approach maximizes the effectiveness of BECoTTA.
Appendix C Additional Quantitative Results
@@ C.1 Results up to round 10
Table 9: Quantitative results of CDS-Easy (balanced weather shifts). We conduct experiments with Cityscapes-to-ACDC benchmarks which contain the weather shifts in the target domains. For a fair comparison, we report both w/o WAD and w/ WAD performance of our models.
@@
BN Stats Adapt (Nado et al., 2020) - 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0
@@
BECoTTA (Ours)-S MoDE 42.9 68.5 35.0 57.2 47.8 50.5 43.0 69.5 35.1 57.2 47.8 50.7 +0.1 0.09M
BECoTTA (Ours)-M MoDE 43.8 68.8 34.9 57.9 49.2 50.9 43.7 68.9 34.8 57.9 49.3 50.9 +0.0 0.63M
BECoTTA (Ours)-L MoDE 43.9 69.1 35.0 58.3 50.2 51.3 44.0 69.1 35.1 58.3 50.2 51.3 +0.0 3.16M
@@
BECoTTA (Ours)-S MoDE 44.1 69.5 40.1 56.8 49.1 51.9 44.0 69.4 40.2 56.9 49.2 51.8 +0.0 0.12M
BECoTTA (Ours)-M MoDE 45.6 70.8 42.6 59.6 50.8 53.9 45.6 70.8 42.6 59.5 50.8 53.9 +0.0 0.77M
BECoTTA (Ours)-L MoDE 45.7 71.4 43.7 59.6 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
@@
BECoTTA (Ours)-S MoDE 42.9 69.5 35.1 57.3 47.8 50.5 43.0 69.5 35.1 57.3 48.8 50.7 +0.1 0.09M
BECoTTA (Ours)-M MoDE 43.7 68.8 34.8 57.9 49.1 50.8 43.7 68.8 34.5 57.9 49.2 50.9 +0.0 0.63M
BECoTTA (Ours)-L MoDE 43.9 69.0 35.0 58.2 50.1 51.3 44.0 69.1 35.0 58.2 50.2 51.3 +0.0 3.16M
@@
BECoTTA (Ours)-S MoDE 44.0 69.4 40.2 56.8 49.2 51.9 44.0 69.4 40.1 56.9 49.1 51.9 +0.0 0.12M
BECoTTA (Ours)-M MoDE 45.6 70.6 42.6 59.6 50.8 53.8 45.6 70.7 42.5 59.5 50.8 53.9 +0.0 0.77M
BECoTTA (Ours)-L MoDE 45.6 71.3 43.7 59.5 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
@@
             + WAD 3.31M 64.6 53.5 62.5 60.1 64.7 53.6 62.5 60.1 64.6 53.5 62.4 60.2 64.5 53.4 62.3 60.1 60.2
C.2 Additional Ablations
@@
More about the routing policy. We also conduct ablation studies on the routing policy for selecting experts within the MoDE layer. We measure the impact of the routing policy in the CDS-Hard scenario under the w/ WAD setting, where [2, 4, 10, 16] hidden dimensions and six experts with $k$=3. The multi-task performance refers to using a fixed assignment per domain, and stochastic routing (Zuo et al., 2021; Wang et al., 2022b) involves random-wise selection. According to  Tab. 13, our chosen top-k routing demonstrates the best performance. This is because the domain-specific router allows for routing that takes input-wise information into consideration.
@@
More about hidden dimension and the number of experts. We include results considering higher hidden dimensions and the number of experts in  Tab. 15, along with the consumption of parameters and memory. The hidden dimension refers to the rank $r$ for each encoder stage block for Segformer (Xie et al., 2021). For instance, [2,4,10,16] means each $r$=2,4,10,16 used at the MoDE layer for four stages Segformer (Xie et al., 2021). [0,0,0,16] denotes that the MoDE layer is used only in the last stage of the encoder. In particular, we predominantly opt for relatively low hidden dimensions and fewer experts, considering the trade-off with efficiency, even though setting a higher hidden dimension generally ensures better performance with more parameters.
C.3 Metrics for Continual Learning

### v3

@@ 4.3 Main Results
Continual Gradual Shifts (CGS). As shown in Tab. 3, we display the first round CGS scenario including four tasks. Even though the target domain is the same setting as CDS-Easy, the overall performances are measured higher since the accessibility of previous domains. Our BECoTTA(L) achieves +5.5%p higher performance than EcoTTA with a similar number of parameters (3.16M). In this case, the input-dependent process of BECoTTA performs well in these blurry scenarios and ultimately shows +13.4%p improvement over the source model.
Classification. In addition to evaluating our method on segmentation tasks, we provide results of CIFAR100-to-CICAR100C classifications to further demonstrate the generalizability of BECoTTA. As shown in Tab. 4, BECoTTA (w/o SDA) consistently outperforms strong baselines, showing -6.82%p and -2.47%p reduction of the average error rate over CoTTA and EcoTTA, respectively. In particular, there are significant improvements over EcoTTA in Contrast (-8.57%p) and Fog (-5.41%p) which have similar attributes to the weather domain. See further results and analyses including computational efficiency in Tab. 12 and Tab. 13.
@@
Zero-shot Domain Generalization (DG). We further demonstrate the versatility of our method through the zero-shot evaluation on four well-known driving datasets: BDD100k, Mapillary, CTAV, and Synthia. We compare the zero-shot performance of models with two different backbones, Deeplab v3+ and Segformer-B2. As shown in Tab. 5, our proposed method continuously outperforms strong baselines, demonstrating the competitive potential of the generalization ability over unseen domains.
@@ 4.4 Analyses and Ablations
Ablation for each element. We observe the variation of the main components: Domain Discriminator (DD), MoDE, and domain-expert synergy loss $\Theta(D;A)$. As shown in Fig. 5 right, relying solely on DD results in only +0.04%p improvement compared with the source only. However, when incorporating the MoDE and $\Theta(D;A)$, there are +5.36%p and +6.12%p improvement respectively, comparing with the source model.
@@
Dependency on SDA. To validate the independence of BECoTTA on SDA, we perform an ablation with various combinations of augmented source domains. For example, D=2 (Source, Night) indicates SDA where the source domain and Night style augmented data (brightness adjusted to be darker), are included. As Tab. 6 represents, our MoDE layer allows a domain-specific adaptation regardless of the relevancy between SDA and target (test-time) domains. For instance, even if SDA only constitutes (Source, Bright) domains, BECoTTA+ responds to the Night target domain effectively, resulting in 8.5%p higher IoU in the night domain than w/o SDA.
(a) Domain Relation
DD MoDE $\Theta(D;A)$ Avg IoU
@@
(b) Ablation of BECoTTA+
Figure 5: Domain Relation and Ablation study. (a): The domain relation between pre-defined SDA dataset and target domains.(b): The ablation study for each element of BECoTTA+. AvgIoU is measured among 10 rounds.
Relation between SDA and $\bm{X_{t}}$. During the CTTA, we generate pseudo domain labels $d$ for each $X_{t}$ (target domains) through DD. These labels are used for initial domain-router assignment to facilitate domain-adaptive routing. Therefore, the relevance between the SDA and the target domain is crucial. In Fig. 5 left, we represent the relationship between pre-defined SDA and $X_{t}$, and our DD faithfully reflects this connection.
Pseudo label analysis. In  Fig. 4, we compare the generated pseudo labels after finishing the ten rounds. Our method exhibits robustness to forgetting in pseudo-label generation compared to other models. For other baselines, there is an erosion of minor labels by the effect of dominant labels (e.g., sky). However, our BECoTTA prevents such occurrences and effectively preserves pseudo labels as each round goes by, and it demonstrates significant efficacy in preserving fine-grained labels.
Table 6: Ablation study for the number of SDA domains.
B-Clear A-Fog A-Night A-Snow B-Overcast Avg
CoTTA (Wang et al., 2022a) 43.3 67.3 34.8 56.9 48.8 50.2
EcoTTA (Song et al., 2023) 41.9 66.1 31.5 55.3 46.2 48.2
BECoTTA (w/o SDA) 43.0 69.5 35.1 57.3 48.8 50.7
D=2 (Source, Night) 44.2 68.9 39.5 56.8 49.4 51.7
D=2 (Source, Bright) 43.9 69.2 38.1 57.7 49.3 51.6
D=3 (Source, Night, Bright) 44.1 69.3 40.4 57.1 49.1 52.0
@@ 5 Conclusion
Acknowledgements
This work was supported by Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIP) (No.2019-0-00075 Artificial Intelligence Graduate School Program (KAIST), No.2022-0-00713), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. RS-2023-00256259).
@@ Appendix A Baselines
In this section, we provide the details of the TTA baselines we use in our main paper. We illustrate the details of other CTTA baselines in Fig. 2.
@@ B.1 Implementation Details
Dataset SDA Target domains
@@ B.2 The details of BECoTTA
Details of w/ & w/o SDA. To ensure a fair comparison, we conduct all experiments in w/ & w/o SDA settings. In the case w/o SDA, it is impossible to collect priors for domain candidates, therefore we adopt a single router $G$, similar to the conventional stochastic routing (Zuo et al., 2021). In settings w/ SDA, as described in the paper, we employ domain-adaptive routing and a domain-experts synergy loss. This approach maximizes the effectiveness of BECoTTA.
B.3 Algorithm of BECoTTA
To clarify our whole CTTA process, we provide the whole pipeline of BECoTTA at Algo. 1. According to the different initialization steps (Line 1 9), BECoTTA can be initialized in various ways including SDA or not.
Algorithm 1 Continual Test-time Adaptation Pipeline
Input: Source domain $X_{s}$, a sequence of target domains $X_{t}=\{X_{t}^{1},X_{t}^{2},\ldots\}$, source model $f$, trainable parts of MoDE $W_{g}^{d}$, $W_{\text{noise}}^{d}$, $W^{\text{down}}$, $W^{\text{up}}$, number of experts $N$, number of domain routers $D$.
1:  # Initialization
2:  if SDA init then
3:     SDA = DomainAugment($X_{s}$)
4:     Update $W_{g}^{d}$, $W_{\text{noise}}^{d}$, $W^{\text{down}}$, $W^{\text{up}}$, Domain Discriminator $DD$ using SDA with $L_{\text{init}}$
5:  else if Source init then
6:     Update $W_{g}^{d}$, $W_{\text{noise}}^{d}$, $W^{\text{down}}$, $W^{\text{up}}$ using Source domain $X_{s}$ with $L_{\text{seg}}$
7:  else
8:     Randomly initialize $W_{g}^{d}$, $W_{\text{noise}}^{d}$, $W^{\text{down}}$, $W^{\text{up}}$
9:  end if
10:  # CTTA
11:  for target domain index $c=1,2,\ldots$ do
12:     for minibatch $x\sim X_{t}^{c}$ do
13:       if SDA init then
14:          $d=DD(x)$
15:       else
16:          $d=\text{Uniform}(0,D)$
17:       end if
18:       for ViT block $\mathbb{B}\sim f$ do
19:          $x=\mathbb{B}(x)$
20:          $h_{d}(x)\leftarrow G_{d}(x),W^{\text{down}},W^{\text{up}}$
21:          $x\leftarrow x+h_{d}(x)$
22:       end for
23:       Update $f$ with $L_{\text{tta}}$
24:     end for
25:  end for
Appendix C Additional Results
@@ C.1 Results up to round 10
Table 9: Quantitative results of CDS-Easy (balanced weather shifts). We conduct experiments with Cityscapes-to-ACDC benchmarks which contain the weather shifts in the target domains. For a fair comparison, we report both w/o SDA and w/ SDA performance of our models.
@@
BN Stats Adapt (Nado et al., 2020a) - 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0
@@
BECoTTA (S) MoDE 42.9 68.5 35.0 57.2 47.8 50.5 43.0 69.5 35.1 57.2 47.8 50.7 +0.1 0.09M
BECoTTA (M) MoDE 43.8 68.8 34.9 57.9 49.2 50.9 43.7 68.9 34.8 57.9 49.3 50.9 +0.0 0.63M
BECoTTA (L) MoDE 43.9 69.1 35.0 58.3 50.2 51.3 44.0 69.1 35.1 58.3 50.2 51.3 +0.0 3.16M
@@
BECoTTA+ (S) MoDE 44.1 69.5 40.1 56.8 49.1 51.9 44.0 69.4 40.2 56.9 49.2 51.8 +0.0 0.12M
BECoTTA+ (M) MoDE 45.6 70.8 42.6 59.6 50.8 53.9 45.6 70.8 42.6 59.5 50.8 53.9 +0.0 0.77M
BECoTTA+ (L) MoDE 45.7 71.4 43.7 59.6 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
@@
BECoTTA (S) MoDE 42.9 69.5 35.1 57.3 47.8 50.5 43.0 69.5 35.1 57.3 48.8 50.7 +0.1 0.09M
BECoTTA (M) MoDE 43.7 68.8 34.8 57.9 49.1 50.8 43.7 68.8 34.5 57.9 49.2 50.9 +0.0 0.63M
BECoTTA (L) MoDE 43.9 69.0 35.0 58.2 50.1 51.3 44.0 69.1 35.0 58.2 50.2 51.3 +0.0 3.16M
@@
BECoTTA+ (S) MoDE 44.0 69.4 40.2 56.8 49.2 51.9 44.0 69.4 40.1 56.9 49.1 51.9 +0.0 0.12M
BECoTTA+ (M) MoDE 45.6 70.6 42.6 59.6 50.8 53.8 45.6 70.7 42.5 59.5 50.8 53.9 +0.0 0.77M
BECoTTA+ (L) MoDE 45.6 71.3 43.7 59.5 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
@@
             + SDA 3.31M 64.6 53.5 62.5 60.1 64.7 53.6 62.5 60.1 64.6 53.5 62.4 60.2 64.5 53.4 62.3 60.1 60.2
Table 12: Classification error rate (%) for CIFAR10-to-CIFAR10C with severity level 5. Results are evaluated on WideResNet-28. * indicates our implemented version performances.
Method Gaus. Shot Impu. Defo. Glas. Moti. Zoom Snow Fros. Fog Brig. Cont. Elas. Pixe. Jpeg Avg. err
Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
tBN (Nado et al., 2020b) 28.6 26.8 37.0 13.2 35.4 14.4 12.6 18.0 18.2 16.0 8.6 13.3 24.0 20.3 27.8 20.9
Single do. TENT (Wang et al., 2020) 25.2 23.8 33.5 12.8 32.3 14.1 11.7 16.4 17.0 14.4 8.4 12.2 22.8 18.0 24.8 19.2
Continual TENT (Wang et al., 2020) 25.2 20.8 29.4 14.4 31.5 15.4 14.2 18.8 17.5 17.3 10.9 14.9 23.6 20.2 25.6 20.0
TTT++ (Liu et al., 2021b) 27.9 25.8 35.8 13.0 34.3 14.2 12.2 17.4 17.6 15.5 8.6 13.1 23.1 19.6 26.6 20.3
SWRNSP (Choi et al., 2022b) 24.6 20.5 29.3 12.4 31.1 13.0 11.3 15.3 14.7 11.7 7.8 9.3 21.5 15.6 20.3 17.2
NOTE (Gong et al., 2022) 30.4 26.7 34.6 13.6 36.3 13.7 13.9 17.2 15.8 15.2 9.1 7.5 24.1 18.4 25.9 20.2
EATA (Niu et al., 2022) 23.8 18.8 27.3 13.9 29.7 16.0 13.3 18.0 16.9 15.7 10.5 12.2 22.9 17.1 23.0 18.6
CoTTA (Wang et al., 2022a) 24.3 21.6 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
CoTTA* 24.6 21.6 26.5 12.1 28.0 13.0 10.9 15.3 14.6 13.6 8.1 12.2 20.0 14.9 19.5 17.0
EcoTTA (k=4) (Song et al., 2023) 23.5 19.0 26.6 11.5 28.1 13.1 10.9 15.2 14.5 13.1 7.8 11.4 20.9 15.4 20.8 16.9
EcoTTA (k=4)* 25.7 21.5 28.4 11.4 31.0 14.1 11.9 16.7 15.3 13.9 8.9 12.4 20.4 16.1 20.7 17.9
BECoTTA (w/o SDA) 22.9 19.1 26.9 10.2 27.5 12.7 10.4 14.7 14.3 12.4 7.2 9.4 20.9 15.2 20.2 16.3
Table 13: Classification error rate (%) for standard CIFAR10-to-CIFAR10C with inference time and memory consumption (MB).
Method Avg. Err. Time (s) Memory (MB)
TENT (Wang et al., 2020) 20.0 1222 118.2
CoTTA (Wang et al., 2022a) 17.0 18877 537.0
EcoTTA* (Song et al., 2023) 17.9 4429 328.3
Ours (Exp10, k4) 16.3 2475 211.6
C.2 Results on Classification Tasks
To validate the versatility of BECoTTA, we conduct additional experiments on classification task scenarios in Tab. 12 and Tab. 4. We also provide the computational inference time/memory efficiency in Tab. 13. All of our BECoTTA are initialized without SDA, which means it ensures a fair comparison with other CTTA baselines. For other baselines’ performances, we borrow performances from Table 19 of the EcoTTA paper.
For the CIFAR10-CIFAR10C task, we adopt the WideResNet-28 backbone. As shown in Tab. 12, our BECoTTA achieves a lower error rate by 4.2%p while reducing inference time by 86.8%p compared to CoTTA, demonstrating that BECoTTA without SDA initialization consistently shows improved performance with remarkable parameter efficiency compared to strong CTTA baselines. Here, we additionally measure the performance of EcoTTA based on the community re-implementation version [^3] https://github.com/Lily-Le/EcoTTA since the official code and checkpoints are not public.
For the CIFAR100-CIFAR100C task, we adopt the WideResNet-40 backbone. While CIFAR100 is a larger scale dataset than CIFAR10, as shown in Tab. 4, BECoTTA outperforms all of the other CTTA baselines in most of the sections.
C.3 Additional Ablations
@@
More about the routing policy. We also conduct ablation studies on the routing policy for selecting experts within the MoDE layer. We measure the impact of the routing policy in the CDS-Hard scenario under the w/ SDA setting, where [2, 4, 10, 16] hidden dimensions and six experts with $k$=3. The multi-task performance refers to using a fixed assignment per domain, and stochastic routing (Zuo et al., 2021; Wang et al., 2022b) involves random-wise selection. According to  Tab. 15, our chosen top-k routing demonstrates the best performance. This is because the domain-specific router allows for routing that takes input-wise information into consideration.
@@
More about hidden dimension. We include results considering higher hidden dimensions and the number of experts in  Tab. 21, along with the consumption of parameters and memory. The hidden dimension refers to the rank $r$ for each encoder stage block for Segformer (Xie et al., 2021). For instance, [2,4,10,16] means each $r$=2,4,10,16 used at the MoDE layer for four stages Segformer (Xie et al., 2021). [0,0,0,16] denotes that the MoDE layer is used only in the last stage of the encoder. In particular, we predominantly opt for relatively low hidden dimensions and fewer experts, considering the trade-off with efficiency, even though setting a higher hidden dimension generally ensures better performance with more parameters.
C.4 Metrics for Continual Learning